ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 92, 1995

Alonso, A. I. and R. Obaya Linear invariant measures for recurrent linear systems

185

Amour, L. Almost periodicity for the nonlinear cubic Schrödinger equation

333

Assani, I., E. Lesigne and D. Rudolph Wiener–Wintner returntimes ergodic theorem

375

Balcar, B. and P. Simon Baire number of the spaces of uniform ultrafilters

263

Berend, D., M. D. Boshernitzan and G. Kolesnik Distribution modulo 1 of some oscillating sequences, II

125

Berkovich, V. G. On the comparison theorem for étale cohomology of non-Archimedean analytic spaces

45

Boshernitzan, M. D. See D. Berend

Caro, Y. Zero-sum subsequences in abelian non-cyclic groups  221

Deshouillers, J-M. and G. A. Freiman On an additive problem of Erdős and Straus, 1

33

Dilworth, S. J. and A. L. Koldobsky The Fourier transform of order statistics with applications to Lorentz spaces

411

Donin, J. and D. Gurevich Some Poisson structures associated to Drinfeld–Jimbo R-matrices and their quantization

23

Drensky, V. New central polynomials for the matrix algebras  235

van den Dries, L. and C. Miller Correction to “On the real exponential field with restricted analytic functions” by Lou van den Dries and Chris Miller, Israel Journal of Mathematics, Vol. 85, 1994, pp. 19–56

427

Freiman, G. A. See J-M. Deshouillers

Ginzburg, D. On the symmetric fourth power L-function of GL_{2}  157
Gitik, M. Some results on the nonstationary ideal  61

Gurevich, D. See J. Donin

Judah, H. and M. Repický No random reals in countable support iterations

349

Kojman, M. and S. Shelah Universal abelian groups  113

Koldobsky, A. L. See S. J. Dilworth Kolesnik, G. See D. Berend Lesigne, E. See I. Assani Livné, R. Motivic orthogonal two-dimensional representations of

Gal(ℚ̄/ℚ) 149

Margolis, S. W. and M. V. Sapir Quasi-identities of finite semigroups and symbolic dynamics

317

Masuoka, A. Semisimple Hopf algebras of dimension 6, 8  361

Miller, C. See L. van den Dries Obaya, R. See A. I. Alonso Repický, M. See H. Judah Rudolph, D. See I. Assani Sapir, M. V. See S. W. Margolis Shelah, S. See M. Kojman Simon, P. See B. Balcar Simon, R. S., S. Spież and H. Toruńczyk The existence of

equilibria in certain games, separation for families of convex functions and a theorem of Borsuk–Ulam type

1

Spież, S. See R. S. Simon

Taskinen, J. Linearization of holomorphic mappings on C(K)-spaces 207

Toruńczyk, H. See R. S. Simon Watanabe, T. The local theta correspondence for unramified unitary

groups

397

Zhang, Q. An example of compact extensions of Kronecker factors and skew products of irrational rotations with finite groups 249